# Aligning Perception, Reasoning, Modeling and Interaction: A Survey on Physical AI

### v3

@@ C.1 Object Recognition
The most fundamental aspect of visual physical perception is the ability to identify objects and determine their spatial relationships within a given scene. In the past decade, the development of convolutional neural networks (CNNs) has made it possible to solve target detection and object classification problems. GPT-4V [44] serves as a milestone that demonstrates robust zero-shot object detection and localization capabilities of MLLMs across diverse visual contexts. Other open-sourced models have also revealed that the introduction of high-quality labeled data can help agents recognize objects at multiple levels of granularity, from basic categories (e.g., "vehicle," "animal") to fine-grained classifications (e.g., "Maserati," "golden retriever") [72, 73, 74]. For more complex scene-level recognition tasks, MLLMs should integrate individual object detections into a coherent understanding of the environment. This involves recognizing scene categories (indoor/outdoor, kitchen/bedroom), understanding typical object arrangements, and identifying anomalous configurations [75, 76, 77, 78]. Recent evaluations [79, 80, 81, 82] have shown that MLLMs have achieved competitive performance on object-level detection capabilities, which lays the foundation for subsequent physical property perception works.
@@ C.2 Spatial Perception
Beyond object recognition, AI systems must understand spatial relationships to build coherent scene representations. This includes both absolute positioning (e.g., "at the center of the whole image") and relative positioning (e.g., "to the left of the white table"). Many recent benchmarks are specifically designed to test spatial understanding, such as VSR and SpatialBench [83, 84, 85, 86, 87], reveal varying capabilities across different MLLMs. Models generally perform well on basic 2D spatial prepositions (above, below, left, right) but struggle with more complex spatial concepts such as 3D spatial reasoning, pixel-level localization and scale relationships. The ability to handle these complex spatial tasks is typically limited by the nature of their training data and architectures. Most models rely on large-scale datasets like COCO [88] or ImageNet [89], which focus on object identification and simple relationships but lack detailed annotations for more sophisticated spatial reasoning tasks. To overcome these limitations, recent research has begun to explore incorporation of spatially aware neural networks [90, 91, 92, 93] and multi-view learning [94, 95, 96, 97] to handle complex spatial relationships.
@@ C.3 Identifying Intrinsic Property
Understanding the physical world from vision requires not only recognizing objects but also inferring their intrinsic properties and dynamic behaviors based on these properties. Intrinsic properties such as mass, viscosity and rigidity are inherent characteristics of objects that remain constant regardless of observational perspectives. Estimating these properties from visual observation alone is particularly challenging for AI as it demands mapping visual features to physical attributes that may not be directly observable but can only be inferred by the laws of physics. Recent studies aim to achieve reliable identification of material such as metal, fabric and plastic [98, 99, 100] and even finer textures (e.g., silk vs. cotton) [101, 102, 103], while also identifying optical properties such as transparency and translucency [104, 105, 106, 107, 108]. Nevertheless, correctly identifying ambiguous cases, novel materials, and lighting variability remain major challenges for frontier models. For mass and weight estimation, models primarily rely on size cues, material-density associations, and category priors, showing stronger performance in relative rather than absolute weight prediction, though systematic biases persist [109, 110, 111, 112, 113, 114]. In terms of rigidity and deformability, MLLMs can classify rigid and flexible objects, detect evidence of deformation, and exploit contextual cues such as force interactions, but their judgments remain heuristic and fragile in complex settings [115, 116, 117, 118, 119].
@@ C.4 Dynamic Estimation
Crucially, these intrinsic properties serve as foundation for understanding how objects behave according to the laws of physics. Knowledge of mass and rigidity, for instance, directly informs whether/how an object might deform under external force, or how it influences motion during collisions. Building on this static perspective, dynamic property perception captures how objects behave and interact over time through contact, constraints, and forces such as support, occlusion, friction, and impact. Unlike intrinsic estimation that answers what an object is in terms of characteristics, dynamic perception addresses how it behaves upon interaction with other objects. Graph-based methods, especially graph neural networks (GNNs), have been central to this effort, with models such as Interaction Networks, Visual Interaction Networks, and the Neural Physics Engine [4, 13, 14] demonstrating how object relations can be captured to predict physical dynamics from visual inputs. More recent frameworks like I-PHYRE [6] further challenge agents to exhibit intuitive physical reasoning, multi-step planning, and in-situ intervention tasks, highlighting the importance of real-time dynamic perception. In addition, benchmarks such as DeepPHY [120] provide systematic evaluation environments that couple intrinsic attributes with dynamic outcomes, ensuring models are tested on both the static and temporal dimensions of physical understanding. By moving from the recognition of intrinsic object properties to the perception of their dynamic relations, AI systems develop a more comprehensive human-like understanding of physical environments.
@@ C.5 Causal and Counterfactual Inference
Beyond perceiving object dynamics from massive aggregated data, causal and counterfactual inference represents the most sophisticated layer of physical reasoning, which focus on elucidating how underlying mechanisms that govern observed phenomena enable reasoning about hypothetical "what if" scenarios without direct observation. Rather than merely describing temporal event sequences, causal reasoning addresses more fundamental questions such as why events occur and how modifications to underlying conditions would alter seemingly irrelevant outcomes. For instance, altering gravitational constants would fundamentally change the falling dynamics of objects, exemplifying the advanced reasoning capabilities required to address such counterfactual queries that transcend pattern matching from observational data. Contemporary methodologies in this domain encompass causal graph modeling [23, 24], intervention-based learning frameworks [121], and hybrid architectures that integrate neural networks with physics engines or symbolic reasoning systems, as demonstrated in PhysORD [25]. Evaluation of these approaches relies on the development of various physical reasoning benchmarks: the Causal3D benchmark [122] provides structured data with corresponding visual representations; the CLEVRER-Humans dataset [123] offers video-based reasoning tasks for causal judgment of physical events with human annotations, addressing the need for diverse event types and natural language descriptions; and the recent PhySense benchmark [124] emphasizes principle-based reasoning by systematically evaluating whether large language models can apply human-like physical principles. By capturing latent causal structures, this layer provides interpretable, generalizable, and mechanistically grounded reasoning capabilities, representing the pinnacle of visual physical understanding.
D Physics Reasoning
As a higher-order cognitive capability that builds upon perceptual foundations, Physics Reasoning operates in the realm of equations, logical inference, and formal problem-solving frameworks. Here, the challenge shifts from handling data complexity to connecting abstract symbolic concepts with the concrete physical phenomena they describe. While perception answers "what is happening", reasoning addresses "why it happens" by applying the structured knowledge embedded in physical laws. In this section, we examine how AI systems perform symbolic physics reasoning from two perspectives: general physics problems (textbook and competition level) and theoretical problems (research level).
@@ D Physics Reasoning
TABLE I: Performance (%) of LLMs on representative text-based and multimodal physics benchmarks. GPT: GPT-4o [28], Claude: Claude-3.7-Sonnet [32], DeepSeek: DeepSeek-R1 [39], Gemini: Gemini-2.5-pro. [33]
@@
UGPhysics [125] 38.66 56.34
PHYBench [126] 6.97 12.87 33.45 45.82
GPQA [127] 50.3 67.4 71.5 84.0
OlympiadBench [128] 39.72 7.34
@@
SeePhys [5] 21.9 34.6 42.2 54.9
PhysReason [129] 29.58 34.07
MMMU [130] 59.4 75.0 81.7
MMMU-Pro [131] 51.9 76.4
@@ D.1 Benchmarking Physics Reasoning
We first systematically review the current landscape of physics reasoning benchmarks from both text-based and multimodal perspectives, shown in Table I.
@@
Representative resources such as PhysicsEval [132] compile problems across multiple education levels and emphasize inference-time techniques, numerical calculation, and robustness under distributional shifts. These benchmarks provide systematic coverage from fundamental mechanics and kinematics to electricity, thermodynamics, and modern physics, enabling evaluation of both single-step calculations and multi-step reasoning chains. Other text-based datasets, including UGPhysics [125], PHYBench [126], GPQA [127], further extend the scope of questions to university-level courses and examinations.
@@
Typical multimodel benchmarks like SeePhys [5] highlight the importance of diagram-dependent physical reasoning by organizing problems across seven domains and twenty-one categories, demonstrating that visual context is often indispensable for accurate solutions. Building on this line, PhysReason [129] expands the paradigm by introducing multimodal diagnostic tasks and stepwise scoring to assess models’ ability to integrate perceptual cues with symbolic reasoning. More recent benchmarks, including ScienceQA [133] for science-oriented diagram comprehension, MMMU and MMMU-Pro [130, 131] for large-scale perception–reasoning evaluation, and SciEval [134] for competition-level and interdisciplinary multimodal science problems, further broden the evaluation landscape by emphasizing visual perception, causal inference, and counterfactual reasoning under dynamically varied conditions. Together, these benchmarks extend the evaluation focus from purely text-based problem solving to scenarios where perception and reasoning must jointly operate in complex visual environments.
@@ D.2 General Physics Reasoning Models
AI systems for general physical reasoning have evolved from simple programs to sophisticated autonomous agents. These systems must tackle complex problems requiring advanced deduction and abstraction of real-world scenarios. Early studies used attention mechanisms to jointly identify salient information in text and diagrams for scientific QA tasks [135]. More recent work in Prompt Engineering and Mixture of Experts (MoE) uses Chain-of-Thought (CoT) prompting to elicit analogical reasoning [136, 137, 138]. Knowledge graph techniques also help by decomposing complex questions into simpler subproblems [139].
A significant leap is the development of agentic systems. A key example is The AI Scientist, which generates research papers from ideation to final manuscript [140]. Its successor, The AI Scientist-V2, was the first AI system to have a paper accepted through peer review [141]. Other agents demonstrate performance equal to human experts in high-level competitions. In particular, Physics Supernova [142] ranked 14th out of 406 contestants at the 2025 International Physics Olympiad (IPhO).
Multi-agent systems have also shown great value for scientific discovery [143] as we demonstrate in Figure 5. For instance, AtomAgents helps design novel metal alloys [144]. The Multimodal Universe project processes vast astronomical datasets and has compiled over 100 TB of data for machine learning research [145]. Visual and multimodal reasoning remains a critical challenge. One successful approach used text descriptions of images to help AI better understand physics diagrams, which won first place at the ICML 2025 AI4Math Challenge [34]. These advancements show a clear trajectory toward automating scientific workflows. However, the field still relies on specialized systems, particularly for tasks with strong visual dependence.
@@ D.3 Theoretical Physics Solvers
The synergy between artificial intelligence and the physical sciences has evolved into two primary paradigms [146, 2]: data-driven scientific discovery and physics-informed learning. These approaches provide the methodological foundation for the advanced reasoning systems discussed previously, each addressing the challenge of integrating physical principles in a distinct way.
Data-driven approaches aim to discover unknown physical laws directly from data. A prominent example is Symbolic Regression (SR), which extracts interpretable mathematical expressions without prior knowledge of the governing equations. This makes SR a powerful tool for generating human-readable models that overcome the "black-box" nature of many neural networks. The primary obstacle for SR is navigating its vast combinatorial search space while ensuring physical plausibility. To this end, recent advancements focus on incorporating hard physical constraints, such as unit consistency, to accelerate the learning process and guide the discovery toward physically meaningful equations [147].
On the other hand, physics-informed learning is applied to problems where the governing equations are already established. This paradigm incorporates known physical laws as constraints within neural architectures. The most notable example is Physics-Informed Neural Networks (PINNs), which embed the governing partial differential equations (PDEs) directly into the neural network’s loss function. This ensures that the model’s solution adheres to fundamental physical principles. This methodology has found widespread application across diverse fields, including the modeling of high-speed and hidden fluid mechanics [148, 149], solving the incompressible Navier-Stokes equations [150], and simulating flow in porous media [151]. Its utility also extends to solid mechanics for problems in elasticity theory and fatigue life prediction [152, 153], as well as to energy systems for solving optimal power flow problems and modeling the complex degradation processes in lithium-ion batteries [154, 155].

### v4

@@ C.1 Object Recognition
The most fundamental aspect of visual physical perception is the ability to identify objects and determine their spatial relationships within a given scene. In the past decade, the development of convolutional neural networks (CNNs) has made it possible to solve target detection and object classification problems. GPT-4V [GPT4V] serves as a milestone that demonstrates robust zero-shot object detection and localization capabilities of MLLMs across diverse visual contexts. Other open-sourced models have also revealed that the introduction of high-quality labeled data can help agents recognize objects at multiple levels of granularity, from basic categories (e.g., "vehicle," "animal") to fine-grained classifications (e.g., "Maserati," "golden retriever") [he2018maskrcnn, liu2024groundingdinomarryingdino, xu2025qwen3omnitechnicalreport]. For more complex scene-level recognition tasks, MLLMs should integrate individual object detections into a coherent understanding of the environment. This involves recognizing scene categories (indoor/outdoor, kitchen/bedroom), understanding typical object arrangements, and identifying anomalous configurations [Song_2024, varghese2025viewinvariantpixelwiseanomalydetection, zhang2024cognition, Ji_2024]. Recent evaluations [MME, Roboflow100-VL, EagleVision, ROD-MLLM] have shown that MLLMs have achieved competitive performance on object-level detection capabilities, which lays the foundation for subsequent physical property perception works.
@@ C.2 Spatial Perception
Beyond object recognition, AI systems must understand spatial relationships to build coherent scene representations. This includes both absolute positioning (e.g., "at the center of the whole image") and relative positioning (e.g., "to the left of the white table"). Many recent benchmarks are specifically designed to test spatial understanding, such as VSR and SpatialBench [SpatialScore, Open3DVQA, yang2025mmsi, wu2025spatialmllmboostingmllmcapabilities, xu2025multi], reveal varying capabilities across different MLLMs. Models generally perform well on basic 2D spatial prepositions (above, below, left, right) but struggle with more complex spatial concepts such as 3D spatial reasoning, pixel-level localization and scale relationships. The ability to handle these complex spatial tasks is typically limited by the nature of their training data and architectures. Most models rely on large-scale datasets like COCO [coco] or ImageNet [Imagenet], which focus on object identification and simple relationships but lack detailed annotations for more sophisticated spatial reasoning tasks. To overcome these limitations, recent research has begun to explore incorporation of spatially aware neural networks [SpIRL, SpaGAN, benou2025tellvisuallyexplainabledeep, cho2024spatially] and multi-view learning [yu2025review, ViewSpatial, zheng2023comprehensive, zhang2024multi] to handle complex spatial relationships.
@@ C.3 Identifying Intrinsic Property
Understanding the physical world from vision requires not only recognizing objects but also inferring their intrinsic properties and dynamic behaviors based on these properties. Intrinsic properties such as mass, viscosity and rigidity are inherent characteristics of objects that remain constant regardless of observational perspectives. Estimating these properties from visual observation alone is particularly challenging for AI as it demands mapping visual features to physical attributes that may not be directly observable but can only be inferred by the laws of physics. Recent studies aim to achieve reliable identification of material such as metal, fabric and plastic [kocsis2024intrinsicimagediffusionindoor, Van_Zuijlen_2021] and even finer textures (e.g., silk vs. cotton) [hu2020fabricsurfacecharacterizationassessment, jang2024fabric], while also identifying optical properties such as transparency and translucency [chen2024practical, weidenbach2024transparency, nagai2025top]. Nevertheless, correctly identifying ambiguous cases, novel materials, and lighting variability remain major challenges for frontier models. For mass and weight estimation, models primarily rely on size cues, material-density associations, and category priors, showing stronger performance in relative rather than absolute weight prediction, though systematic biases persist [Vision-Based, afridi2024analyzing, kalus2024simulating, M_ller_2024, nath2024mass]. In terms of rigidity and deformability, MLLMs can classify rigid and flexible objects, detect evidence of deformation, and exploit contextual cues such as force interactions, but their judgments remain heuristic and fragile in complex settings [PhysID, mustafa2025forcemapping, song2025moda, PaNDaS].
@@ C.4 Dynamic Estimation
Crucially, these intrinsic properties serve as foundation for understanding how objects behave according to the laws of physics. Knowledge of mass and rigidity, for instance, directly informs whether/how an object might deform under external force, or how it influences motion during collisions. Building on this static perspective, dynamic property perception captures how objects behave and interact over time through contact, constraints, and forces such as support, occlusion, friction, and impact. Unlike intrinsic estimation that answers what an object is in terms of characteristics, dynamic perception addresses how it behaves upon interaction with other objects. Graph-based methods, especially graph neural networks (GNNs), have been central to this effort, with models such as Interaction Networks, Visual Interaction Networks, and the Neural Physics Engine [VIN, NPE] demonstrating how object relations can be captured to predict physical dynamics from visual inputs. More recent frameworks like I-PHYRE [I-PHYRE] further challenge agents to exhibit intuitive physical reasoning, multi-step planning, and in-situ intervention tasks, highlighting the importance of real-time dynamic perception. In addition, benchmarks such as DeepPHY [DeepPHY] provide systematic evaluation environments that couple intrinsic attributes with dynamic outcomes, ensuring models are tested on both the static and temporal dimensions of physical understanding. By moving from the recognition of intrinsic object properties to the perception of their dynamic relations, AI systems develop a more comprehensive human-like understanding of physical environments.
@@ C.5 Causal and Counterfactual Inference
Beyond perceiving object dynamics from massive aggregated data, causal and counterfactual inference represents the most sophisticated layer of physical reasoning, which focus on elucidating how underlying mechanisms that govern observed phenomena enable reasoning about hypothetical "what if" scenarios without direct observation. Rather than merely describing temporal event sequences, causal reasoning addresses more fundamental questions such as why events occur and how modifications to underlying conditions would alter seemingly irrelevant outcomes. For instance, altering gravitational constants would fundamentally change the falling dynamics of objects, exemplifying the advanced reasoning capabilities required to address such counterfactual queries that transcend pattern matching from observational data. Contemporary methodologies in this domain encompass causal graph modeling [causalmodeling], intervention-based learning frameworks [ke2020causalmodels], and hybrid architectures that integrate neural networks with physics engines or symbolic reasoning systems, as demonstrated in PhysORD [PhysORD2024]. Evaluation of these approaches relies on the development of various physical reasoning benchmarks: the Causal3D benchmark [causal3D] provides structured data with corresponding visual representations; the CLEVRER-Humans dataset [clevrerhumans] offers video-based reasoning tasks for causal judgment of physical events with human annotations, addressing the need for diverse event types and natural language descriptions; and the recent PhySense benchmark [physense] emphasizes principle-based reasoning by systematically evaluating whether large language models can apply human-like physical principles. By capturing latent causal structures, this layer provides interpretable, generalizable, and mechanistically grounded reasoning capabilities, representing the pinnacle of visual physical understanding.
D Physics Reasoning: From Observation to Explanation
Physics reasoning represents the critical transition from perception to explanation, where systems move beyond identifying physical properties to understanding why phenomena occur. Building directly upon perceptual foundations, reasoning operates through equations and theoretical principles that encode physical laws. While perception answers "what is happening" by extracting observable properties, reasoning addresses "why it happens" by connecting abstract concepts with concrete phenomena through causal relationships governed by natural laws. This capability bridges observational understanding with predictive modeling, enabling systems to not merely recognize patterns but derive explanations grounded in physical principles. We examine how AI systems perform physics reasoning across two domains of increasing sophistication: general physics problems spanning textbook exercises to competition-level challenges, and theoretical problems requiring research-level mathematical derivation with conceptual synthesis. This progression reflects the deepening cognitive demands from applying established principles to discovering novel theoretical insights, ultimately preparing systems for the modeling stage where learned physical laws enable forward prediction and counterfactual simulation.
@@ D Physics Reasoning: From Observation to Explanation
TABLE II: Performance (%) of LLMs on representative text-based and multimodal physics benchmarks. GPT: GPT-4o [gpt4o], Claude: Claude-3.7-Sonnet [Claude-3.7-Sonnet], DeepSeek: DeepSeek-R1 [guo2025deepseek], Gemini: Gemini-2.5-pro. [Gemini-pro]
@@
UGPhysics [ugphysics] 38.66 56.34
PHYBench [qiu2025phybench] 6.97 12.87 33.45 45.82
GPQA [GPQA] 50.3 67.4 71.5 84.0
OlympiadBench [olympiadbench] 39.72 7.34
@@
SeePhys [seephys] 21.9 34.6 42.2 54.9
PhysReason [physreason] 29.58 34.07
MMMU [mmmu] 59.4 75.0 81.7
MMMU-Pro [mmmupro] 51.9 76.4
@@ D.1 Benchmarking Physics Reasoning
We first systematically review the current landscape of physics reasoning benchmarks from both text-based and multimodal perspectives, shown in Table II.
@@
Representative resources such as PhysicsEval [physicseval] compile problems across multiple education levels and emphasize inference-time techniques, numerical calculation, and robustness under distributional shifts. These benchmarks provide systematic coverage from fundamental mechanics and kinematics to electricity, thermodynamics, and modern physics, enabling evaluation of both single-step calculations and multi-step reasoning chains. Other text-based datasets, including UGPhysics [ugphysics], PHYBench [qiu2025phybench], GPQA [GPQA], further extend the scope of questions to university-level courses and examinations.
@@
Typical multimodel benchmarks like SeePhys [seephys] highlight the importance of diagram-dependent physical reasoning by organizing problems across seven domains and twenty-one categories, demonstrating that visual context is often indispensable for accurate solutions. Building on this line, PhysReason [physreason] expands the paradigm by introducing multimodal diagnostic tasks and stepwise scoring to assess models’ ability to integrate perceptual cues with symbolic reasoning. More recent benchmarks, including ScienceQA [scienceqa] for science-oriented diagram comprehension, MME-SCI and MMMU series [mme_sci, mmmu, mmmupro] for large-scale perception–reasoning evaluation, and SciEval [scieval] for competition-level and interdisciplinary multimodal science problems, further broden the evaluation landscape by emphasizing visual perception, causal inference, and counterfactual reasoning under dynamically varied conditions. Together, these benchmarks extend the evaluation focus from purely text-based problem solving to scenarios where perception and reasoning must jointly operate in complex visual environments.
@@ D.2 General Physics Reasoning Models
AI systems for general physical reasoning have evolved from simple programs to sophisticated autonomous agents. These systems must tackle complex problems requiring advanced deduction and abstraction of real-world scenarios. Early studies used attention mechanisms to jointly identify salient information in text and diagrams for scientific QA tasks [GmezPrez2020ISAAQMC]. More recent work in Prompt Engineering and Mixture of Experts (MoE) uses CoT prompting to elicit analogical reasoning [Addala2024StepsAAA, Jaiswal2024ImprovingPR]. Decomposing problems or reasoning processes into simpler atomic components has also been shown to improve reasoning performance, as demonstrated by instruction fine-tuning methods [xiang2024atomthink, kuang2025atomic] and knowledge graph-based approaches [Addala2024KnowledgeGAE].
A significant leap is the development of agentic systems. A key example is The AI Scientist, which generates research papers from ideation to final manuscript [lu2024aiscientist]. Its successor, The AI Scientist-V2, was the first AI system to have a paper accepted through peer review [yamada2024aiscientistv2]. Other agents demonstrate performance equal to human experts in high-level competitions. In particular, Physics Supernova [PhysicsSupernova] ranked 14th out of 406 contestants at the 2025 International Physics Olympiad (IPhO).
Multi-agent systems have also shown great value for scientific discovery [CollectiveIntelligence] as we demonstrate in Figure 5. For instance, AtomAgents helps design novel metal alloys [ghafarollahi2024atomagents]. The Multimodal Universe project processes vast astronomical datasets and has compiled over 100 TB of data for machine learning research [qu2024multimodal]. Visual and multimodal reasoning remains a critical challenge. One successful approach used text descriptions of images to help AI better understand physics diagrams, which won first place at the ICML 2025 AI4Math Challenge [Liang2025MultimodalRF]. These advancements show a clear trajectory toward automating scientific workflows. However, the field still relies on specialized systems, particularly for tasks with strong visual dependence.
@@ D.3 Theoretical Physics Solvers
The synergy between artificial intelligence and the physical sciences has evolved into two primary paradigms [carleo2019_ml_physics_survey, AIMeetPhy_survey_2024]: data-driven scientific discovery and physics-informed learning. These approaches provide the methodological foundation for the advanced reasoning systems discussed previously, each addressing the challenge of integrating physical principles in a distinct way.
Data-driven approaches aim to discover unknown physical laws directly from data. A prominent example is Symbolic Regression (SR), which extracts interpretable mathematical expressions without prior knowledge of the governing equations. This makes SR a powerful tool for generating human-readable models that overcome the "black-box" nature of many neural networks. The primary obstacle for SR is navigating its vast combinatorial search space while ensuring physical plausibility. To this end, recent advancements focus on incorporating hard physical constraints, such as unit consistency, to accelerate the learning process and guide the discovery toward physically meaningful equations [zhang2024_interpretable_turbulence_sr].
On the other hand, physics-informed learning is applied to problems where the governing equations are already established. This paradigm incorporates known physical laws as constraints within neural architectures. The most notable example is Physics-Informed Neural Networks (PINNs), which embed the governing partial differential equations (PDEs) directly into the neural network’s loss function. This ensures that the model’s solution adheres to fundamental physical principles. This methodology has found widespread application across diverse fields, including the modeling of high-speed and hidden fluid mechanics [mao2020_pinn_high_speed_flow, raissi2020_hidden_fluid_mechanics], solving the incompressible Navier-Stokes equations [jin2021_nsfnets], and simulating flow in porous media [almajid2022_pinn_porous_media]. Its utility also extends to solid mechanics for problems in elasticity theory and fatigue life prediction [vahab2021_pinn_biharmonic_elasticity, zhou2023_pinn_fatigue_life], as well as to energy systems for solving optimal power flow problems and modeling the complex degradation processes in lithium-ion batteries [nellikkath2022_pinn_ac_opf, wang2024_pinn_battery_degradation].